# Matrix Rosenthal and Concentration Inequalities for Markov Chains with Applications in Statistical Learning

### v1

@@ 5 Conclusions
In this work, we have studied probability bounds for the spectral norm of sums of dependent random matrices. First, we have proved a Rosenthal-Burkholder inequality for matrix martingales using the extrapolation method (good $\lambda$ inequality) (Hitczenko, 1990, 1994; Pinelis, 1994) and matrix Bennett inequality (Tropp, 2011; Kroshnin and Suvorikova, 2024). Furthermore, combining this martingale inequality with the Poisson equation approach (Durmus et al., 2023), we have derived matrix Rosenthal inequalities for Markov chains under both the uniform geometric ergodicity assumption and the geometric $V$-ergodicity assumption. By Markov inequality, these moment inequalities can be further translated into concentration inequalities including Hoeffding and Bernstein inequalities. Through more careful analysis, we have also established dimension-free versions of the above inequalities.
Compared with previous results based on the León-Perron argument (León and Perron, 2004), our results yield a leading term that matches the Markov chain CLT, rather than relying on a variance proxy ($t_{\operatorname{mix}}\|{\mathbb{E}}_{\pi}[{\bm{F}}^{2}(Z_{0})]\|$ or $t_{\operatorname{mix}}\|\|{\bm{F}}\|\|^{2}$). And our results do not require the assumption of a non-zero absolute spectral gap. This expands the applicability to a broader range of problems, such as policy evaluation in reinforcement learning (Duan and Wainwright, 2022) and PCA (Kumar and Sarkar, 2023) with Markovian data. In contrast to prior works relying on the multi-matrix Golden-Thompson inequality (Sutter et al., 2017; Garg et al., 2018), our results can be extended to dimension-free versions, enabling applications to infinite-dimensional settings. Moreover, our approach is not restricted to the uniform geometric ergodicity assumption, so we can derive the results under more general ergodicity assumptions, and in some cases we do not require the functions to be bounded. The advantages of our results over previous works are demonstrated in the covariance estimation and PCA problems on Markovian data.
One direction for future work is to explore whether we can obtain residual terms with sharp dependence on $t_{\operatorname{mix}}$. This might require more sophisticated tools from matrix analysis, beyond the properties of self-adjoint matrices related to the Löwner order. Another promising direction is to extend our results to time-inhomogeneous functions while retaining the advantages mentioned above. For time-inhomogeneous functions, the standard Poisson equation approach no longer works, and more advanced tools from Markov chain theory would be necessary.
@@ Appendix A Omitted Proof for the Rosenthal-Burkholder Inequality for Matrix Martingales
$\left\{{\sup_{k\in{\mathbb{N}}}\|{{\bm{M}}}_{k}\|>\lambda}\right\}=\left\{{\tau_{1}<\infty}\right\},\quad\left\{{\sup_{k\in{\mathbb{N}}}\|{{\bm{M}}}_{k}\|>\beta\lambda}\right\}=\left\{{\tau_{2}<\infty}\right\},\quad\left\{{\left\|{\left\langle{\bar{\bm{M}}}\right\rangle_{\infty}}\right\|^{1/2}\leq\delta_{1}\lambda}\right\}=\left\{{\tau_{3}=\infty}\right\}.$ (18)
@@
which is the desired conclusion. Here, we explain some details in the derivation of Eqn (19). In the first inequality, we used the fact that $\langle\bar{{\bm{M}}}\rangle_{\infty}\preccurlyeq\langle{\bm{M}}\rangle_{\infty}$ almost surely. In the second inequality, we used Eqn. (18). In the third inequality, we used the fact that on the event $\{\tau_{2}<\infty,\tau_{3}=\infty,\sup_{k\in{\mathbb{N}}}\|{\bm{X}}_{k}\|\leq\delta_{2}\lambda\}$, it holds that $\sup_{k<\tau_{1}}\left\|{{{\bm{M}}}_{k}}\right\|=\sup_{k<\tau_{1}}\left\|{\bar{{\bm{M}}}_{k}}\right\|<\lambda$ and $\left\|{\bar{{\bm{M}}}_{(\tau_{2}\wedge\tau_{3})-1}}\right\|>\left\|{{{\bm{M}}}_{\tau_{2}}}\right\|-\left\|{{\bm{X}}_{\tau_{2}}}\right\|>(\beta-\delta_{2})\lambda$. And in the last inequality, we used the matrix Bennett inequality (Lemma A.1) with error term $(\beta-1-\delta_{2})\lambda$ parameter $U=\delta_{2}\lambda$ and $\sigma=\delta_{1}\lambda$.
@@
If we further assume that there exists a ${\bm{\Upsilon}}\in{\mathbb{H}}^{d}_{+}$ such that $\langle{\bm{M}}\rangle_{\infty}\preccurlyeq{\bm{\Upsilon}}$ almost surely, then the third inequality also holds if we replace $d$ with $e\cdot r({\bm{\Upsilon}}\wedge(\delta_{1}^{2}\lambda^{2}))$ by the dimension-free matrix Bennett inequality (Lemma A.1). ∎
A.2 Proof of Lemma 4.1
@@
A.3 Remaining Proof of Theorem 2.1

### v2

@@ 5 Conclusions
In this work, we have studied probability bounds for the spectral norm of sums of dependent random matrices. First, we have proved a Rosenthal-Burkholder inequality for matrix martingales (Theorem 2.1) using the extrapolation method (good $\lambda$ inequality) (Hitczenko, 1990, 1994; Pinelis, 1994) and the matrix Bennett inequality (Tropp, 2011; Kroshnin and Suvorikova, 2024). By discretization, we have shown a sharp Burkholder-Davis-Gundy inequality for continuous matrix local martingales (Theorem 2.2) as a corollary of Theorem 2.1. Furthermore, combining this martingale inequality with the Poisson equation approach (Durmus et al., 2023), we have derived matrix Rosenthal inequalities for Markov chains under both the uniform geometric ergodicity assumption and the geometric $V$-ergodicity assumption in Section 2.2. By the Markov inequality, these moment inequalities can be further translated into concentration inequalities including Hoeffding and Bernstein inequalities. Through more careful analysis, we have also established dimension-free versions of the above inequalities.
Compared with previous works on matrix concentration inequalities for Markov chains, our results yield a sharp leading term that matches the Markov chain CLT, rather than relying on a sub-optimal variance proxy. Our results do not require the assumption of a non-zero absolute spectral gap or $\psi$-mixing condition. This expands its applicability to a broader range of problems, such as policy evaluation in reinforcement learning (Duan and Wainwright, 2022) and PCA (Kumar and Sarkar, 2023) with Markovian data. And our results do not need the assumption of the boundedness of functions under the geometric ergodicity assumption. Moreover, our results can be extended to dimension-free versions, enabling applications to infinite-dimensional settings. The advantages of our results over previous works are demonstrated in the covariance estimation and PCA problems on Markovian data. And as mentioned in this paper, our results can also be extended to more general geometrically ergodic Markov chains with respect to the Wasserstein semimetric, as well as to geometrically ergodic continuous Markov processes.
One direction for future work is to explore whether we can obtain residual terms with sharp dependence on $t_{\operatorname{mix}}$. This might require more sophisticated tools from matrix analysis, beyond the properties of self-adjoint matrices related to the Löwner order. Another promising direction is to extend our results to time-inhomogeneous functions while retaining the advantages mentioned above. For time-inhomogeneous functions, the standard Poisson equation approach no longer works, and more advanced tools from Markov chain theory would be necessary. Finally, an interesting question is: do certain form of universality-based probabilistic inequalities hold for martingales and ergodic Markov chains? The universality-based inequalities may yield tighter error bounds in some applications, and we would leave the question for future work.
@@ Appendix A Omitted Proof for the Rosenthal-Burkholder Inequality for Matrix Martingales
[…]
@@
which is the desired conclusion. Here, we explain some details in the derivation of Eqn (21). In the first inequality, we used the fact that $\langle\bar{{\bm{M}}}\rangle_{\infty}\preccurlyeq\langle{\bm{M}}\rangle_{\infty}$ almost surely. In the second inequality, we used Eqn. (20). In the third inequality, we used the fact that on the event $\{\tau_{2}<\infty,\tau_{3}=\infty,\sup_{k\in{\mathbb{N}}}\|{\bm{X}}_{k}\|\leq\delta_{2}\lambda\}$, it holds that $\sup_{k<\tau_{1}}\left\|{{{\bm{M}}}_{k}}\right\|=\sup_{k<\tau_{1}}\left\|{\bar{{\bm{M}}}_{k}}\right\|<\lambda$ and $\left\|{\bar{{\bm{M}}}_{(\tau_{2}\wedge\tau_{3})-1}}\right\|>\left\|{{{\bm{M}}}_{\tau_{2}}}\right\|-\left\|{{\bm{X}}_{\tau_{2}}}\right\|>(\beta-\delta_{2})\lambda$. And in the last inequality, we used the matrix Bennett inequality (Lemma A.1) with error term $\varepsilon=(\beta-1-\delta_{2})\lambda$, parameter $U=\delta_{2}\lambda$, and $\|\langle\hat{{\bm{M}}}_{k}\rangle_{\infty}\|\leq\delta_{1}^{2}\lambda^{2}$ given ${\mathcal{F}}_{\tau_{1}}$.
@@
If we further assume that there exists a ${\bm{\Upsilon}}\in{\mathbb{H}}^{d}_{+}$ such that $\langle{\bm{M}}\rangle_{\infty}\preccurlyeq{\bm{\Upsilon}}$ almost surely, then the fourth inequality also holds if we replace $d$ with $e\cdot r({\bm{\Upsilon}}\wedge(\delta_{1}^{2}\lambda^{2}))$ by the dimension-free matrix Bennett inequality (Lemma A.1). ∎
A.2 Proof of Theorem 2.2
Proof.
We first consider $\tau=\infty$ and the bounded case: $({\bm{M}}_{t})_{t\geq 0}$ is a continuous martingale satisfying $\sup_{t\geq 0}\left\|{{\bm{M}}_{t}}\right\|\leq K,\left\|{\left\langle{{\bm{M}}}\right\rangle_{\infty}}\right\|\leq K$ for some constant $K>0$. For a fixed discretization level $n\in{\mathbb{N}}$, we define ${\bm{M}}_{k}^{(n)}={\bm{M}}_{\frac{k}{2^{n}}}$ for any $k\in{\mathbb{N}}$ (${\bm{M}}_{-1}^{(n)}:=\bm{0}_{d{\times}d}$). Apply Theorem 2.1 to $({\bm{M}}_{k}^{(n)})_{k\in{\mathbb{N}}}$, we have
$\left\|{\sup_{k\in{\mathbb{N}}}\|{\bm{M}}_{\frac{k}{2^{n}}}\|}\right\|_{p}$ (22)
$\leq$ $C_{R,1}\sqrt{p\vee\log d}\left\|{\left\|{\sum_{k=0}^{\infty}{\mathbb{E}}\left[{({\bm{M}}_{\frac{k}{2^{n}}}-{\bm{M}}_{\frac{k-1}{2^{n}}})^{2}\big|{{\mathscr{F}}_{\frac{k-1}{2^{n}}}}}\right]}\right\|^{1/2}}\right\|_{p}+C_{R,2}(p\vee\log d)\left\|{\sup_{k\in{\mathbb{N}}}\|{\bm{M}}_{\frac{k}{2^{n}}}-{\bm{M}}_{\frac{k-1}{2^{n}}}\|}\right\|_{p}.$
Now we take the limit of each part of the above inequality as $n\to\infty$. By Beppo Levi’s Theorem, the left hand side converges to $\left\|{\sup_{t\geq 0}\|{\bm{M}}_{t}\|}\right\|_{p}$. By dominated convergence theorem, the second term of the right hand side converges to 0. It suffices to prove $\lim_{n\to\infty}\sum_{k=0}^{\infty}{\mathbb{E}}\left[{({\bm{M}}_{\frac{k}{2^{n}}}-{\bm{M}}_{\frac{k-1}{2^{n}}})^{2}\big|{{\mathscr{F}}_{\frac{k-1}{2^{n}}}}}\right]=\left\langle{{\bm{M}}}\right\rangle_{\infty}$ in $L^{\frac{p}{2}}$. Here the convergence of matrices is defined entry-wise.
Lemma A.2.
[…]
Proof.
${\mathbb{E}}\left\|{\sum_{k=0}^{\infty}({\bm{M}}_{\frac{k}{2^{n}}}-{\bm{M}}_{\frac{k-1}{2^{n}}})^{2}}\right\|^{2}$ $\leq{\mathbb{E}}\left({\operatorname{Tr}\left({{\sum_{k=0}^{\infty}({\bm{M}}_{\frac{k}{2^{n}}}-{\bm{M}}_{\frac{k-1}{2^{n}}})^{2}}}\right)}\right)^{2}$ (24)
$={\mathbb{E}}\left({\sum_{i,j=1}^{d}{\sum_{k=0}^{\infty}\left({[{\bm{M}}_{\frac{k}{2^{n}}}]_{ij}-[{\bm{M}}_{\frac{k-1}{2^{n}}}]_{ij}}\right)^{2}}}\right)^{2}$
$\leq d^{2}\sum_{i,j=1}^{d}{\mathbb{E}}\left({{\sum_{k=0}^{\infty}\left({[{\bm{M}}_{\frac{k}{2^{n}}}]_{ij}-[{\bm{M}}_{\frac{k-1}{2^{n}}}]_{i,j}}\right)^{2}}}\right)^{2}.$
By [Karatzas and Shreve, 1988, Lemma 1.5.9], each term in the last summation is upper bounded by $48K^{4}$. ∎
Lemma A.3.
$\lim_{n\to\infty}{\mathbb{E}}\left\|{\sum_{k=0}^{\infty}({\bm{M}}_{\frac{k}{2^{n}}}-{\bm{M}}_{\frac{k-1}{2^{n}}})^{4}}\right\|=0$ (25)
Proof.
${\mathbb{E}}\left\|{\sum_{k=0}^{\infty}({\bm{M}}_{\frac{k}{2^{n}}}-{\bm{M}}_{\frac{k-1}{2^{n}}})^{4}}\right\|$ $\leq{\mathbb{E}}\left\|{\sum_{k=0}^{\infty}\left({\sup_{k\in{\mathbb{N}}}\|{\bm{M}}_{\frac{k}{2^{n}}}-{\bm{M}}_{\frac{k-1}{2^{n}}}\|^{2}}\right)({\bm{M}}_{\frac{k}{2^{n}}}-{\bm{M}}_{\frac{k-1}{2^{n}}})^{2}}\right\|$ (26)
$={\mathbb{E}}\left({\sup_{k\in{\mathbb{N}}}\|{\bm{M}}_{\frac{k}{2^{n}}}-{\bm{M}}_{\frac{k-1}{2^{n}}}\|^{2}}\right)\left\|{\sum_{k=0}^{\infty}({\bm{M}}_{\frac{k}{2^{n}}}-{\bm{M}}_{\frac{k-1}{2^{n}}})^{2}}\right\|$
$\leq\left({{\mathbb{E}}\sup_{k\in{\mathbb{N}}}\|{\bm{M}}_{\frac{k}{2^{n}}}-{\bm{M}}_{\frac{k-1}{2^{n}}}\|^{4}}\right)^{{1}/{2}}\left({{\mathbb{E}}\left\|{\sum_{k=0}^{\infty}({\bm{M}}_{\frac{k}{2^{n}}}-{\bm{M}}_{\frac{k-1}{2^{n}}})^{2}}\right\|^{2}}\right)^{{1}/{2}}$
$\leq\sqrt{48}d^{2}K^{2}\left({{\mathbb{E}}\sup_{k\in{\mathbb{N}}}\|{\bm{M}}_{\frac{k}{2^{n}}}-{\bm{M}}_{\frac{k-1}{2^{n}}}\|^{4}}\right)^{{1}/{2}}.$
In the last step, we use the upper bound in Lemma A.2. The expectation in the last step converges to 0 by dominated convergence theorem. ∎
Lemma A.4.
$\lim_{n\to\infty}\sum_{k=0}^{\infty}{\mathbb{E}}\left[{({\bm{M}}_{\frac{k}{2^{n}}}-{\bm{M}}_{\frac{k-1}{2^{n}}})^{2}\big|{{\mathscr{F}}_{\frac{k-1}{2^{n}}}}}\right]=\left\langle{{\bm{M}}}\right\rangle_{\infty}\ {\rm in}\ L^{2}.$ (27)
Proof.
Note that $(({\mathbb{E}}[({\bm{M}}_{\frac{k}{2^{n}}}-{\bm{M}}_{\frac{k-1}{2^{n}}})^{2}|{{\mathscr{F}}_{\frac{k-1}{2^{n}}}}]-(\left\langle{{\bm{M}}}\right\rangle_{\frac{k}{2^{n}}}-\left\langle{{\bm{M}}}\right\rangle_{\frac{k-1}{2^{n}}})))_{k\in{\mathbb{N}}}$ is a martingale difference sequence. Thus,
${\mathbb{E}}\left\|{\sum_{k=0}^{\infty}\left({{\mathbb{E}}\left[{({\bm{M}}_{\frac{k}{2^{n}}}-{\bm{M}}_{\frac{k-1}{2^{n}}})^{2}\big|{{\mathscr{F}}_{\frac{k-1}{2^{n}}}}}\right]-(\left\langle{{\bm{M}}}\right\rangle_{\frac{k}{2^{n}}}-\left\langle{{\bm{M}}}\right\rangle_{\frac{k-1}{2^{n}}})}\right)}\right\|_{F}^{2}$ (28)
$=$ $\sum_{k=0}^{\infty}{\mathbb{E}}\left\|{\left({{\mathbb{E}}\left[{({\bm{M}}_{\frac{k}{2^{n}}}-{\bm{M}}_{\frac{k-1}{2^{n}}})^{2}\big|{{\mathscr{F}}_{\frac{k-1}{2^{n}}}}}\right]-(\left\langle{{\bm{M}}}\right\rangle_{\frac{k}{2^{n}}}-\left\langle{{\bm{M}}}\right\rangle_{\frac{k-1}{2^{n}}})}\right)}\right\|_{F}^{2}$
$\leq$ $2\sum_{k=0}^{\infty}{\mathbb{E}}\left\|{{\mathbb{E}}\left[{({\bm{M}}_{\frac{k}{2^{n}}}-{\bm{M}}_{\frac{k-1}{2^{n}}})^{2}\big|{{\mathscr{F}}_{\frac{k-1}{2^{n}}}}}\right]}\right\|_{F}^{2}+2\sum_{k=0}^{\infty}{\mathbb{E}}\left\|{\left\langle{{\bm{M}}}\right\rangle_{\frac{k}{2^{n}}}-\left\langle{{\bm{M}}}\right\rangle_{\frac{k-1}{2^{n}}}}\right\|_{F}^{2}$
$\leq$ $2\sum_{k=0}^{\infty}{\mathbb{E}}\left\|{({\bm{M}}_{\frac{k}{2^{n}}}-{\bm{M}}_{\frac{k-1}{2^{n}}})^{2}}\right\|_{F}^{2}+2\sum_{k=0}^{\infty}{\mathbb{E}}\left\|{\left\langle{{\bm{M}}}\right\rangle_{\frac{k}{2^{n}}}-\left\langle{{\bm{M}}}\right\rangle_{\frac{k-1}{2^{n}}}}\right\|\operatorname{Tr}\left({\left\langle{{\bm{M}}}\right\rangle_{\frac{k}{2^{n}}}-\left\langle{{\bm{M}}}\right\rangle_{\frac{k-1}{2^{n}}}}\right)$
$\leq$ $2\sum_{k=0}^{\infty}{\mathbb{E}}\operatorname{Tr}\left({{\bm{M}}_{\frac{k}{2^{n}}}-{\bm{M}}_{\frac{k-1}{2^{n}}}}\right)^{4}+2\sum_{k=0}^{\infty}{\mathbb{E}}\left({\sup_{k\in{\mathbb{N}}}\left\|{\left\langle{{\bm{M}}}\right\rangle_{\frac{k}{2^{n}}}-\left\langle{{\bm{M}}}\right\rangle_{\frac{k-1}{2^{n}}}}\right\|}\right)\operatorname{Tr}\left({\left\langle{{\bm{M}}}\right\rangle_{\frac{k}{2^{n}}}-\left\langle{{\bm{M}}}\right\rangle_{\frac{k-1}{2^{n}}}}\right)$
$=$ $2{\mathbb{E}}\operatorname{Tr}\sum_{k=0}^{\infty}({\bm{M}}_{\frac{k}{2^{n}}}-{\bm{M}}_{\frac{k-1}{2^{n}}})^{4}+2{\mathbb{E}}\left({\sup_{k\in{\mathbb{N}}}\left\|{\left\langle{{\bm{M}}}\right\rangle_{\frac{k}{2^{n}}}-\left\langle{{\bm{M}}}\right\rangle_{\frac{k-1}{2^{n}}}}\right\|}\right)\operatorname{Tr}(\left\langle{{\bm{M}}}\right\rangle_{\infty}),$
where the first term converges to 0 by Lemma A.3, and the second term converges to 0 by the dominated convergence theorem. ∎
Lemma A.5.
$\left({\left\|{\sum_{k=0}^{\infty}{\mathbb{E}}\left[{({\bm{M}}_{\frac{k}{2^{n}}}-{\bm{M}}_{\frac{k-1}{2^{n}}})^{2}\big|{{\mathscr{F}}_{\frac{k-1}{2^{n}}}}}\right]}\right\|^{\frac{p}{2}}}\right)_{n\in{\mathbb{N}}}$ (29)
is uniformly integrable.
Proof.
Take $q>\frac{p}{2}$. It suffices to prove
$\left({{\mathbb{E}}\left\|{\sum_{k=0}^{\infty}{\mathbb{E}}\left[{({\bm{M}}_{\frac{k}{2^{n}}}-{\bm{M}}_{\frac{k-1}{2^{n}}})^{2}\big|{{\mathscr{F}}_{\frac{k-1}{2^{n}}}}}\right]}\right\|^{q}}\right)_{n\in{\mathbb{N}}}$ (30)
is uniformly bounded.
${\mathbb{E}}\left\|{\sum_{k=0}^{\infty}{\mathbb{E}}\left[{({\bm{M}}_{\frac{k}{2^{n}}}-{\bm{M}}_{\frac{k-1}{2^{n}}})^{2}\big|{{\mathscr{F}}_{\frac{k-1}{2^{n}}}}}\right]}\right\|^{q}$ $\leq{\mathbb{E}}\left({\operatorname{Tr}\left({\sum_{k=0}^{\infty}{\mathbb{E}}\left[{({\bm{M}}_{\frac{k}{2^{n}}}-{\bm{M}}_{\frac{k-1}{2^{n}}})^{2}\big|{{\mathscr{F}}_{\frac{k-1}{2^{n}}}}}\right]}\right)}\right)^{q}$ (31)
$={\mathbb{E}}\left({\sum_{i,j=1}^{d}\sum_{k=0}^{\infty}{\mathbb{E}}\left[{([{\bm{M}}_{\frac{k}{2^{n}}}]_{ij}-[{\bm{M}}_{\frac{k-1}{2^{n}}}]_{ij})^{2}\big|{{\mathscr{F}}_{\frac{k-1}{2^{n}}}}}\right]}\right)^{q}$
$\leq d^{2(q-1)}\sum_{i,j=1}^{d}{\mathbb{E}}\left({\sum_{k=0}^{\infty}{\mathbb{E}}\left[{\left({[{\bm{M}}_{\frac{k}{2^{n}}}]_{ij}-[{\bm{M}}_{\frac{k-1}{2^{n}}}]_{ij}}\right)^{2}\big|{{\mathscr{F}}_{\frac{k-1}{2^{n}}}}}\right]}\right)^{q}.$
Apply scalar Rosenthal-Burkholder inequality [Burkholder, 1973] to $([{\bm{M}}_{k}^{(n)}]_{ij})_{k\in{\mathbb{N}}}$:
${\mathbb{E}}\left({\sum_{k=0}^{\infty}{\mathbb{E}}\left[{\left({[{\bm{M}}_{\frac{k}{2^{n}}}]_{ij}-[{\bm{M}}_{\frac{k-1}{2^{n}}}]_{ij}}\right)^{2}\big|{{\mathscr{F}}_{\frac{k-1}{2^{n}}}}}\right]}\right)^{q}\leq C_{q}K^{2q}.$ (32)
Thus
$\left({{\mathbb{E}}\left\|{\sum_{k=0}^{\infty}{\mathbb{E}}\left[{({\bm{M}}_{\frac{k}{2^{n}}}-{\bm{M}}_{\frac{k-1}{2^{n}}})^{2}\big|{{\mathscr{F}}_{\frac{k-1}{2^{n}}}}}\right]}\right\|^{q}}\right)_{n\in{\mathbb{N}}}$ (33)
is uniformly bounded by $d^{2q}C_{q}K^{2q}$. ∎
By Lemma A.3 and Lemma A.4, $\lim_{n\to\infty}\sum_{k=1}^{2^{n}}{\mathbb{E}}\left[{({\bm{M}}_{\frac{k}{2^{n}}T}-{\bm{M}}_{\frac{k-1}{2^{n}}T})^{2}\big|{{\mathscr{F}}_{\frac{k-1}{2^{n}}T}}}\right]=\left\langle{{\bm{M}}}\right\rangle_{T}$ in $L^{\frac{p}{2}}$.
Now suppose that $({\bm{M}}_{t})$ is a continuous local martingale that is not necessarily bounded. We use the standard localization technique. Define a sequence of stopping times $\tau_{m}={\rm inf}\{t\geq 0:\left\|{{\bm{M}}_{t}}\right\|\geq m\text{ or }\left\|{\left\langle{{\bm{M}}}\right\rangle_{t}}\right\|\geq m\}$. Then ${\bm{M}}_{t}^{(m)}:={\bm{M}}_{t\wedge\tau\wedge\tau_{m}}$ is a proper martingale. By the previous results, we have
$\left\|{\sup_{t\geq 0}\|{\bm{M}}_{t\wedge\tau\wedge\tau_{m}}\|}\right\|_{p}\leq C_{R,1}\sqrt{p\vee\log d}\Big\|\left\|{\left\langle{{\bm{M}}}\right\rangle_{\tau\wedge\tau_{m}}}\right\|^{1/2}\Big\|_{p}\leq C_{R,1}\sqrt{p\vee\log d}\Big\|\left\|{\left\langle{{\bm{M}}}\right\rangle_{\tau}}\right\|^{1/2}\Big\|_{p}.$ (34)
Let $m\to\infty$ and we finish the proof. ∎
A.3 Proof of Lemma 4.1
@@
A.4 Remaining Proof of Theorem 2.1